# Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models

### v3

@@ 4.2 Identifying Detrimental Priors
We employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images:
$\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.6. We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}_{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.3 and Fig.4. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors. Further discussion can be found in Appendix. C.3
@@
Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$
@@
VCD 85.30 84.59 81.75 82.06 84.74 82.80
ICD 84.81 83.44 80.95 81.59 83.10 80.52
IBD 85.36 84.28 74.82 77.87 83.07 80.23
DeFG 86.68 85.71 84.35 84.50 84.04 81.49
CICD 87.85 87.77 84.00 84.89 87.54 86.37
@@
VCD 80.81 80.84 77.22 78.60 82.95 81.17
ICD 81.44 80.53 76.82 78.48 81.26 78.87
IBD 81.68 81.13 69.35 74.31 82.20 79.59
DeFG 82.96 82.46 79.25 80.46 83.02 80.57
CICD 84.83 85.18 81.01 82.53 85.94 84.90
@@
VCD 77.27 78.19 74.03 76.39 80.37 78.96
ICD 78.07 77.98 73.19 75.91 79.03 77.05
IBD 78.26 78.52 66.60 72.63 79.68 77.36
DeFG 79.40 79.69 75.19 77.52 80.42 78.24
CICD 78.60 80.30 75.55 78.63 81.94 81.37
@@
4.3 Cross-Images Contrastive Decoding
After completing the above preparation, we integrate the proposed method into the contrastive decoding process. Specifically, the model performs regular decoding when essential priors are detected and applies contrastive decoding to suppress detrimental priors. In the case of detrimental priors, smaller distributional differences imply a stronger detrimental effect. Accordingly, we define the contrastive coefficient $\alpha$ in Eq. 3 as: $\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$. Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,x,y_{<t})=$ (6)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 3.
@@ 5 Experiment
Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
5.1 Experiment Settings
Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
Benchmarks
We evaluate the performance of CICD across 4 widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks. These include two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
Evaluated LVLMs
We evaluate the generalizability of CICD across 6 LVLMs from three model families: InstructBLIP Dai et al. (2023), the LLaVA series (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)), and the Qwen series (Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025)). Unless stated otherwise, all models are implemented at the 7B scale.
Baselines
We compare CICD with 4 recent training-free methods that specifically target language priors: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2024b), and DeFG Zhang et al. (2025). All of these methods construct negative contexts based on the original image. In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
5.2 Implementation Details
We use sampling decoding for next-token prediction with default settings. Following Leng et al. (2024), we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. All experiments are conducted on an NVIDIA RTX 3090 GPU.
5.3 Main Results
Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
@@ 5.3 Main Results
Regular 55.7 17.4 52.60 58.6 18.0 52.99 9.9 7.3 30.73
VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7.0 29.70
IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
DeFG 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
Table 4: Results on the COCO subset of DetailCaps. The maximum generation length is 512. Cs, Ci, and CAP represent CHAIRs, CHAIRi, and CAPTURE respectively.
Results on POPE
As shown in Tab. 1, our method achieves the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. The Popular setting uses frequent objects for negative samples, increasing susceptibility to language priors. The strong performance gains of our method in this setting highlight its effectiveness in mitigating such priors.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
@@
LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87
CICD 37.6 10.4 62.7 88.22 86.29 82.76
Qwen2-VL-Instruct Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 50.6 9.4 73.6 90.67 88.07 84.53
Qwen2.5-VL-Instruct Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 39.4 9.6 68.1 89.95 88.00 84.75
Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA).
Results on CHAIR
The results in Tab. 2 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
Results on AMBER
As illustrated in Tab. 3, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
Results on DetailCaps
Tab. 4 presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
5.4 Effectiveness on More LVLMs
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. 5. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. We also test our method on these LVLMs by DetailCaps and the results are shown in Tab. 6. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
Further experiments will be discussed in Sec. B.
6 Conclusion
In this paper, we mitigate the influence of language priors in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
The simplicity and efficiency of CICD offer strong practical potential, enabling seamless integration with other hallucination mitigation methods to further enhance LVLM performance. This integration forms a key direction for our future work.
7 Limitation
Addtional Image
While the additional image is necessary, there is no strict requirement on its degree of difference from the original. Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images. Furthermore, our experiments demonstrate that the effectiveness of hallucination mitigation is not strongly correlated with image similarity. The experimental results demonstrate that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method. This flexibility significantly lowers the barrier for real-world deployment.
Twice Forward Propagation
Appendix A Detailed Related Work
A.1 Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) Dai et al. (2023); Liu et al. (2024c); Bai et al. (2023b); Zhu et al. (2024a) consist of a visual encoder Radford et al. (2021); Dosovitskiy et al. (2021) and a Large Language Model (LLM) backbone Touvron et al. (2023); Chiang et al. (2023). The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning Liu et al. (2023), LVLMs acquire the ability to perform various multimodal tasks such as image captioning Ye et al. (2025), visual question answering Yu et al. (2024), and visual grounding Lu et al. (2024). Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
A.2 Hallucination in LVLMs
Originally, hallucinations Ji et al. (2023); Liu et al. (2024b) were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al. (2023c); Wang et al. (2023), requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al. (2025); Li et al. (2025), attention misallocation Huang et al. (2024); Yin et al. (2025); Kang et al. (2025), and language priors Leng et al. (2024); Zhu et al. (2024b); Wang et al. (2024c). Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
A.3 Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors Zhibo et al. (2023); Wu et al. (2022). Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding Li et al. (2023a); Chuang et al. (2024), achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image Leng et al. (2024), perturbing the instruction Wang et al. (2024c), augmenting the vision input Zhu et al. (2024b), or performing cross-modal conversion Zhang et al. (2025). During decoding, the two logit distributions are contrasted to eliminate language priors.
While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features. Through a comprehensive analysis of language priors, we identify their vision-independent nature. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
Appendix B Detailed Experiment Settings
B.1 Benchmarks

### v5

@@ 3.3 Dynamic Language Prior Suppression
To demonstrate the effectiveness of JS divergence in determining context relevance, we plot the logarithm of the JS divergence between distributions generated with different images at each decoding step. Figure 5 shows the results. Clearly, image-relevant tokens such as “_people”, “_white”, and “_red” exhibit higher divergence values (often above 4-4 when taking $\log_{10}(d(\bm{v},\bm{v}^{\prime}))$), while function words such as “_of”, “_be” and “_the” have lower JS divergence values (often below 4-4 in $\log_{10}(d(\bm{v},\bm{v}^{\prime}))$). We further randomly select 500 images from the MSCOCO dataset, captioned by LLaVA-1.5, and compute the JS divergence between logit distributions produced by different images for each next-token prediction. The statistical results are visualized in Figure 6. We observe that the JS divergence values also fall into two distinct ranges: $\log_{10}(d(\bm{v},\bm{v}^{\prime}))>-4$ and $\log_{10}(d(\bm{v},\bm{v}^{\prime}))\leq-4$. These results further confirm that JS divergence can effectively distinguish between image-relevant and image-irrelevant contexts.
Based on the above analysis, CICD enables dynamic suppression of language priors. Specifically, CICD performs regular decoding at steps where image-irrelevant contexts are detected, while applying contrastive decoding with different images to suppress language priors in image-relevant contexts. A hyperparameter $\gamma$ is used to determine whether the current decoding step is image-irrelevant. Formally, the final logit in CICD is calculated as
$\mathrm{logit}(y_{t})$ $=\left\{\begin{array}[]{ll}\mathrm{logit}_{\theta}(y_{t}\ |\ \bm{v},\mathbf{x},\mathbf{y}_{<t}),&\textrm{if }d(\bm{v},\bm{v}^{\prime})\leq\gamma\\ \mathrm{logit}_{\textrm{CD}}(y_{t}\ |\ \bm{v},\mathbf{x},\mathbf{y}_{<t}),&\textrm{if }d(\bm{v},\bm{v}^{\prime})>\gamma\end{array}\right$. (4)
where $\mathrm{logit}_{\textrm{CD}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})$ is defined as
$\mathrm{logit}_{\textrm{CD}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})=$ $(1+\alpha)\textrm{logit}_{\bm{\theta}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})$ (5)
$-\alpha\textrm{logit}_{\bm{\theta}}(y_{t}|\bm{v}^{\prime},\mathbf{x},\mathbf{y}_{<t}),$
where $\alpha$ is the contrastive coefficient. Unlike previous approaches such as Leng et al. (2024), $\alpha$ is not treated as a hyperparameter in our method. Instead, it is dynamically set to $1-\log_{10}(d(\bm{v},\bm{v}^{\prime}))$. This is based on the intuition that detrimental priors with higher similarity are more difficult to remove and therefore require stronger suppression. To ensure numerical stability, we clip $\alpha$ to the interval $[1,3]$.
@@
[…]
@@
VCD Leng et al. (2024) 85.30 84.59 81.75 82.06 84.74 82.80
ICD Wang et al. (2024b) 84.81 83.44 80.95 81.59 83.10 80.52
PAI Liu et al. (2024d) 84.56 83.33 79.54 80.27 84.14 81.77
IBD Zhu et al. (2025a) 85.36 84.28 74.82 77.87 83.07 80.23
DeGF Zhang et al. (2025) 86.68 85.71 84.35 84.50 84.04 81.49
CICD (ours) 87.85 87.77 84.00 84.89 87.54 86.37
@@
VCD Leng et al. (2024) 80.81 80.84 77.22 78.60 82.95 81.17
ICD Wang et al. (2024b) 81.44 80.53 76.82 78.48 81.26 78.87
PAI Liu et al. (2024d) 80.27 79.72 75.04 77.05 83.06 80.89
IBD Zhu et al. (2025a) 81.68 81.13 69.35 74.31 82.20 79.59
DeGF Zhang et al. (2025) 82.96 82.46 79.25 80.46 83.02 80.57
CICD (ours) 84.83 85.18 81.01 82.53 85.94 84.90
@@
VCD Leng et al. (2024) 77.27 78.19 74.03 76.39 80.37 78.96
ICD Wang et al. (2024b) 78.07 77.98 73.19 75.91 79.03 77.05
PAI Liu et al. (2024d) 77.02 77.30 72.14 75.15 80.62 78.76
IBD Zhu et al. (2025a) 78.26 78.52 66.60 72.63 79.68 77.36
DeGF Zhang et al. (2025) 79.40 79.69 75.19 77.52 80.42 78.24
CICD (ours) 78.60 80.30 75.55 78.63 81.94 81.37
@@ 4 Experiment
4.1 Setup
Benchmarks.
We evaluate our CICD on 4 widely adopted multimodal hallucination benchmarks: DetailCaps Ye et al. (2025): A fine-grained image captioning benchmark. We evaluate on a 700-image MSCOCO subset. POPE Li et al. (2023b): A hallucination evaluation benchmark with 9,000 questions and 1,500 images from MSCOCO Lin et al. (2014), A-OKVQA Schwenk et al. (2022), and GQA Hudson and Manning (2019), covering three negative sampling settings (Random, Popular, and Adversarial). We also evaluate our method on the CHAIR Rohrbach et al. (2018) and AMBER Wang et al. (2023a) benchmarks. Due to space constraints, we leave the results in Appendix D.
LVLMs.
We use 6 representative models to study the effectiveness of CICD, which are InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024b), LLaVA-Next Liu et al. (2024c), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024a), and Qwen2.5-VL-Instruct Bai et al. (2025) respectively. The parameter counts of the above models range from 7B to 8B unless otherwise stated.
Baselines.
We compare CICD with the following representative training-free methods: VCD Leng et al. (2024): a method that introduces noise into the input image during contrastive decoding to suppress language priors. ICD Wang et al. (2024b): a method that also uses contrastive decoding but adopts adversarial instruction to construct negative logits. IBD Zhu et al. (2025a): a method that increases the influence of image information during decoding through strengthening the model’s attention to visual features. PAI Liu et al. (2024d): a method that removes the visual input from the original input and retains the pure textual context as the negative context, while simultaneously enhancing the model’s attention to visual tokens in the original context. DeGF Zhang et al. (2025): a method that converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations. The implementation details of baselines are provided in the Appendix B.
Implementation Details.
We use sampling decoding for next-token prediction with default settings. Following Leng et al. (2024), we incorporate adaptive plausibility constraints into the decoding process. For image retrieval, we use the MSCOCO 2014 validation set as $\mathcal{D}_{\textrm{img}}$, which contains 40,504 images. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Unless otherwise noted, we set the threshold $\gamma=-4$ in all our experiments.
4.2 Main Results
Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
[…]
Regular 55.7 17.4 52.60 58.6 18.0 52.99 9.9 7.3 30.73
VCD Leng et al. (2024) 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
ICD Wang et al. (2024b) 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7.0 29.70
PAI Liu et al. (2024d) 39.4 11.8 53.49 62.6 18.3 53.27 4.6 3.3 29.54
IBD Zhu et al. (2025a) 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
DeGF Zhang et al. (2025) 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
CICD (ours) 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
Table 2: Results on the COCO subset of DetailCaps. Cs, Ci, and CAP denote CHAIRs, CHAIRi, and CAPTURE respectively.
Results on POPE.
Table 1 presents the results of all baseline methods alongside our proposed method. Our approach achieves the highest F1 scores across all sampling settings and backbone LVLMs. These results suggest that CICD effectively mitigates hallucinations and promotes the use of faithful visual information in LVLMs. Notably, under the Popular setting, which uses frequent objects as negative samples and thus increases the likelihood of over-reliance on language priors. Nevertheless, our CICD method still achieves significant improvements. This highlights its effectiveness in reducing dependency on language priors.
Results on DetailCaps.
Table 2 presents the results on the MSCOCO subset of DetailCaps. We primarily use CAPTURE as the main evaluation metric, which assesses caption quality through both lexical matching and semantic alignment by measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our CICD method achieves the best performance across all three LVLMs, demonstrating strong image captioning capabilities and confirming its generalizability across different datasets. These results suggest that CICD is effective in mitigating the over-reliance on language priors.
4.3 Comparison of Retrieval Methods
[…]
Regular 54.6 16.4 72.6 82.25 78.99 77.13
[…]
Table 3: Comparison of CICD with different contrastive input construction methods on the CHAIR and POPE benchmarks. Ran, Pop, and Adv denote the Random, Popular, and Adversarial subsets, respectively.
[…]
We further analyze the similarity between the original and contrastive visual inputs by computing the overlap ratio of image-related words in their descriptions within the COCO dataset. We find that the overlap ratios for random sampling and image retrieval are only 0.24% and 0.11%, respectively, relative to the original inputs, which are significantly lower than those of other methods, as shown in Figure 7. These findings suggest that using a visually distinct image is effective for contrastive decoding. Moreover, random sampling can achieve performance comparable to image retrieval, provided that the dataset $\mathcal{D}_{\textrm{img}}$ contains sufficiently diverse images.
4.4 Analysis
Method CHAIR64 CHAIR512
$downarrow Cs\downarrow Ci\downarrow$
PAI 20.0 6.2 41.2 10.4
PAI (w/ DI) 18.0 5.0 40.2 11.2
VCD 25.0 8.3 59.8 17.8
VCD (w/ DI) 22.6 7.4 45.0 12.1
Table 4: CHAIR64 and CHAIR512 represent results on the CHAIR benchmark with maximum generation lengths of 64 and 512, respectively. (w/ DI) indicates the method with our proposed distinct contrastive inputs.
Distorted Contrastive Distributions.
We analyze the effect of distorted distributions resulting from contrastive visual inputs. As previously discussed, methods such as VCD Leng et al. (2024) and PAI Liu et al. (2024d) introduce a training–inference gap due to the use of noisy or absent visual inputs. To examine whether these distorted distributions negatively impact the performance of contrastive decoding, we directly apply our contrastive input construction method to both VCD and PAI. The results, shown in Table 4, indicate that the performance of both methods improves when using our construction approach. This confirms that distorted contrastive distributions can indeed degrade the effectiveness of contrastive decoding.
[…]
LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87 38.0 10.4 61.06
CICD 37.6 10.4 62.7 88.22 86.29 82.76 36.2 8.5 62.19
Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43 51.3 10.6 60.04
CICD 50.6 9.4 73.6 90.67 88.07 84.53 48.4 9.9 61.79
Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12 42.2 11.7 62.62
CICD 39.4 9.6 68.1 89.95 88.00 84.75 43.8 10.9 65.07
Table 5: Results on additional LVLMs. CHAIR512 denotes results on the CHAIR benchmark with a maximum generation length of 512. Ran., Pop., and Adv. denote the Random, Popular, and Adversarial subsets, respectively.
Figure 7: Overlap ratio of image-related words between original and negative contexts in contrastive decoding, based on 100 randomly selected COCO images.
Figure 8: Correlation between visual information overlap ratio and hallucination behavior.
Visual Information Overlap.
We analyze the visual information overlap between the original and contrastive contexts across different contrastive decoding methods by computing the overlap ratio of image-related words in their descriptions using the MSCOCO dataset. Implementation details can be found in Appendix C. Figure 7 presents the overlap ratios for various methods. The results show that VCD, ICD, and DeGF exhibit substantial overlap with the original input, indicating that their contrastive visual inputs retain a significant amount of shared visual content. In contrast, our method yields a nearly negligible overlap, suggesting it provides more distinct contrastive contexts. The information overlap between the two contexts in contrastive decoding confuses the model, amplifying the advantages of suboptimal tokens and thereby inducing hallucinatory behavior. Furthermore, we analyze the association between overlap ratios and hallucination behaviors based on the DeGF, which has balanced hallucination rates and overlap ratios. We remove samples without hallucinations and without visual information overlap. Although the statistical results in Figure 8 do not show a strong correlation ($p=0.076$), there is still a positive association ($\rho=0.29$). In short, CICD achieves significantly lower hallucination rates compared to VCD, ICD, and DeGF, demonstrating the superiority of low visual overlap in CICD.
Method CHAIR64 CHAIR512
$downarrow Cs\downarrow Ci\downarrow$
Selective Removal of Language Priors
CICD ($\gamma=-4$) 18.0 6.1 43.8 11.7
CICD (w/ DeGF) 21.8 6.9 56.8 16.4
Complete Removal of Language Priors
CICD ($\gamma=$-$\infty$) 19.6 6.1 44.4 11.7
CICD (w/ VCD) 22.6 7.4 45.0 12.1
Table 6: Comparison among different contrastive decoding methods. CICD (w/ VCD) is equivalent to VCD (w/ DI).
Excessive Suppression of Language Priors.
While contrastive decoding is effective in suppressing language priors, it also raises concerns about excessive suppression, which may negatively affect the coherence of generated text. To investigate this issue, we set $\gamma$ to -$\infty$ to mitigate all priors and further replace Eq. 4 with other contrastive decoding methods to evaluate the superiority of dynamic language prior suppression. As shown in Table 6, we observe that using cross-image consistency to control the extent of language prior suppression is highly effective, leading to reduced hallucination. Moreover, preserving essential priors is crucial, as the performance of CICD degrades when $\gamma$ is set to -$\infty$. VCD removes all language priors while DeGF augments the retained language priors, both of which yield suboptimal performance. In particular, DeGF’s augmentation of retained priors can harm model performance when prior issues are more severe, such as with longer contexts. These results indicate that the dynamic language prior suppression method can effectively mitigate language priors without harming the performance of LVLMs. The impact of the choice of $\gamma$ will be discussed in Appendix D.6.
4.5 Effectiveness of CICD on More LVLMs
To comprehensively demonstrate the effectiveness and generalizability of our CICD method, we conduct additional experiments using three of the most recent open-source LVLMs. The results, presented in Table 5, show that our method consistently achieves significant performance improvements across all experiments. CICD effectively reduces hallucinations in both generative and discriminative tasks by mitigating the influence of language priors. These results further confirm the transferability of CICD to a variety of LVLMs.
5 Conclusion
In this paper, we first reveal that language priors are intrinsic to LVLMs and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD), a training-free method that effectively suppresses language priors without compromising the model’s performance. Despite its simplicity and efficiency, CICD demonstrates strong practical potential and can be seamlessly integrated with other hallucination mitigation methods to further enhance LVLM performance.
Acknowledgments
This work is supported by Zhongguancun Academy Project No.20240103.
Appendix A Limitation
Addtional Image
Although the additional image is indispensable, its acquisition is straightforward. Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images. We demonstrate that random selection and retrieval-based methods achieve comparable low visual information overlap ratios when selecting images. The experimental results also show that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method. This flexibility significantly lowers the barrier for real-world deployment.
Twice Forward Propagation
@@ Appendix A Limitation
Table 8: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by (100CHAIR+F1)/2(100-CHAIR+F1)/2.
Appendix B Implementation Details for Baselines
The implementation details for the reproduction of baselines are as follows:
  • VCD Leng et al. (2024): We added 500 steps of Gaussian noise to the images to construct the negative visual input for contrastive decoding. The hyperparameters in VCD are set to $\alpha=1,\beta=0.1$.
  • ICD Wang et al. (2024b): We use the sentence "You are a confused objects detector to provide a fuzzy overview or impression of the image." as the disturbance instruction for contrastive decoding. The hyperparameters for contrastive decoding are set according to VCD.
  • IBD Zhu et al. (2025a): We follow the practice in the original paper, setting $\epsilon=2$, $\alpha=1.5\times 10^{4}$, and $\beta=0.1$ in our experiments.
  • PAI Liu et al. (2024d): We set the visual attention augmentation coefficient $\alpha$ to 0.5 and enable increased visual attention starting from layer 2 (counting from 0). The hyperparameter $\gamma$ is set to 1.1. PAI employs adaptive plausibility constraints with $\beta=0.1$.
  • DeGF Zhang et al. (2025) We employ LLaVA-v1.5-7B to generate captions for each image and use Stable Diffusion v1-5 to generate images from the captions. The parameter $\gamma$ is set to 0.1, and $\beta$ is set to 0.1 for the image captioning task and 0.25 for the visual question answering task.
Note: The same parameters may have different meanings across methods. For specific details, please refer to the original papers.
Appendix C Implementation Details for Visual Information Overlap
First, we apply several contrastive decoding methods, each constructing its own negative context, and use them to prompt LVLMs to answer user queries. For comparison, we obtain standard responses via regular decoding. We then use QwenVL-Max-0125 to extract image-related words from the generated responses. Inspired by CoT Wei et al. (2022), we guide the model through a multi-stage extraction process: it first identifies all visually relevant words without access to the image, and then selects those strongly associated with the image content, which we define as relevant words.
Finally, we compute the overlap ratio between the relevant words generated from negative contexts and those obtained through regular decoding. This ratio quantifies how much essential visual content is preserved in responses generated from negative contexts, serving as an indicator of retained visual information. The prompt used to guide QwenVL-Max is shown in Figure 10.
@@ Appendix C Implementation Details for Visual Information Overlap
LLaVA-1.5-13B Regular 58.8 17.0 73.4 82.94 81.46 78.49
CICD 44.4 12.9 76.0 87.92 85.68 80.99
INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
CICD 44.0 11.6 65.7 87.52 83.60 78.96
Appendix D More Experimental Results
D.1 Benchmarks